# Randomized block-Krylov subspace methods for low-rank approximation of matrix functions

### v1

@@ 3.3 Probabilistic bounds
If $\bm{\Omega}$ is a standard Gaussian matrix, $\bm{\Omega}_{k}$ and $\bm{\Omega}_{n\setminus k}$ are as defined as in (2), and $\mathcal{E}_{\bm{\Omega}}(s;f)$ and $\mathcal{E}(s;f)$ are as defined in Eqs. 10 and 16, then
@@
    for any $u,t\geq 0$, with probability at least $1-e^{-(u-2)/4}-\sqrt{\pi k}\left(\frac{t}{e}\right)^{-(\ell-k+1)/2}$,
    $\mathcal{E}_{\bm{\Omega}}(s;f)\leq\frac{utk}{\ell-k+1}\mathcal{E}(s;f);$
@@
(i): For any polynomial $p\in\mathbb{P}_{s-1}$ by [40, Proposition 8.6] we have with probability at least $1-e^{-(u-2)/4}-\sqrt{\pi k}\left(\frac{t}{e}\right)^{-(\ell-k+1)/2}$
@@
The inequality is respected if we minimize both sides over all polynomials.
(ii): This is proven in an identical fashion utilizing the expectation bound in [40, Proposition 8.6]. ∎
@@ 3.4 Error bounds for Krylov aware low-rank approximation
    with probability at least $1-e^{-(u-2)/4}-\sqrt{\pi r}\left(\frac{t}{e}\right)^{-(\ell-k+1)/2}$,
@@
    $\sqrt{\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}+\frac{5utk}{\ell-k+1}\mathcal{E}(s;f)};$
@@
We conclude this section by commenting on the three terms appearing in the bounds in Theorem 3.5. The dependence on $4\sqrt{\ell s}\inf\limits_{p\in\mathbb{P}_{2r+1}}\|f(x)-p(x)\|_{L^{\infty}([\lambda_{\min},\lambda_{\max}])}$ captures the inherent cost of approximating a quadratic form with $f(\bm{A})$ using the Lanczos method, as in done in Line 3 of Algorithm 4. A similar term would also arise in an analysis of the Randomized SVD when matvecs are computed approximately using a Krylov subspace method. The $\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}$ term is due to the fact that our error can never be below the optimal rank $k$ approximation error. Finally, $\mathcal{E}(s;f)$ tells us that $\bm{Q}_{s}$ is a good orthonormal basis for low-rank approximation if there is a polynomial of degree at most $s-1$ that is large on the eigenvalues $\lambda_{1},\ldots,\lambda_{k}$ (which correspond to the top subspace of $f(\bm{A})$) and is small on the eigenvalues $\lambda_{k+1},\ldots,\lambda_{n}$.
As discussed in Section 3.2, one possible choice of polynomial would be an approximation to $f$, in which case $\mathcal{E}(s;f)$ would be close to $\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}$. In this case, ignoring the polynomial approximation term, we would recover a bound almost identical to the standard Randomized SVD error bound. In particular, when matvecs with $f(\bm{A})$ are implemented exactly, Randomized SVD can be shown to have expected error [21]:
@@
However, other possibly polynomials can also be chosen, so we might have $\mathcal{E}(s;f)\ll\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}$. In the next section, we give an example involving the matrix exponential to better illustrate why this is often the case.
@@ 3.5 Example bounds for the matrix exponential
One matrix function for which it is often desirable to obtain a low-rank approximation is the matrix exponential, $\exp(\bm{A})$. This task arises in tasks ranging from network analysis [28], to quantum thermodynamics [6, 10], to solving PDEs [32]. In this section, we do a deeper dive into our main bounds for Krylov aware low-rank approximation when applied to the matrix exponential to better understand how they compare to bounds obtainable using the more naive approach of directly combining the randomized SVD with a black-box matrix-vector product approximation algorithm for $\exp(\bm{A})$. For conciseness of exposition, we focus on expectation bounds, and defer details of proofs to Appendix A.
We first demonstrates that, as discussed in the previous section, by choosing $p$ in $\mathcal{E}(s;\exp(x))$ to be a polynomial approximation to $\exp(x)$, we can recover the bounds of the randomized SVD [21, Theorem 10.5], up to small factors accounting for the fact that $\exp(x)$ is not a polynomial.
@@
Note that as $\frac{\sqrt{\ell s}\gamma_{1,n}^{2r+2}}{2^{4r+1}(2r+2)!}\|\exp(\bm{A})\|_{2}\to 0$ and $\frac{\gamma_{1,n}^{s}}{s!}\to 0$, we recover the classical bound for the randomized SVD [21, Theorem 10.5]. Furthermore, these two terms converge superexponentially to 00 as $s,r\to+\infty$. However, due to the $\frac{5k}{\ell-k-1}$, the above bound only implies a subexponential convergence to the error of the optimal rank $k$ approximation error, $\|\exp(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}$, in terms of the number of required matrix-vector products with $\bm{A}$, which scales as $\ell(r+s)$.
@@
Consider the setting of Theorem 3.5 with $f(x)=\exp(x)$. Define $\gamma_{i,j}=\lambda_{i}-\lambda_{j}$ to be the gap between the $i^{\text{th}}$ and $j^{\text{th}}$ largest eigenvalues. Then, if $\ell-k\geq 2$ and $s\geq 2+\frac{\gamma_{1,k}}{\log\left(\frac{\gamma_{1,n}}{\gamma_{k,n}}\right)}$
@@
$\hskip 50.00008pt+\sqrt{1+\exp(2\gamma_{k,n})\cdot 2^{-2(s-2)\sqrt{\min\left\{1,2\frac{\gamma_{k,k+1}}{\gamma_{k+1,n}}\right\}}}\frac{80k}{\ell-k-1}}\|\exp(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}.$

### v2

@@ 3.3 Probabilistic bounds
[…]
@@
    for any $\delta\in(0,1)$ and $\ell\geq k$, with probability at least $1-\delta$,
    $\mathcal{E}_{\bm{\Omega}}(s;f)\leq C_{\delta,k,\ell}\mathcal{E}(s;f),$
    where $C_{\delta,k,\ell}=2e\left(2\log(2/\delta)+1\right)\left(\frac{2\sqrt{\pi k}}{\delta}\right)^{\frac{2}{\ell-k+1}}\frac{k}{\ell-k+1}$;
@@
(i): For any polynomial $p\in\mathbb{P}_{s-1}$ by [tropp2023randomized, Proposition 8.6] we have with probability at least $1-e^{-(u-2)/4}-\sqrt{\pi k}\left(\frac{t}{e}\right)^{-(\ell-k+1)/2}$
@@
The inequality is respected if we minimize both sides over all polynomials. Solving $e^{(u-2)/2}\leq\delta/2$ and $\sqrt{\pi k}\left(\frac{t}{e}\right)^{-(\ell-k+1)/2}\leq\delta/2$ for $u$ and $t$ yields the desired result.
(ii): This is proven in an identical fashion utilizing the expectation bound in [tropp2023randomized, Proposition 8.6]. ∎
@@ 3.4 Error bounds for Krylov aware low-rank approximation
    with probability at least $1-\delta$,
@@
    $\sqrt{\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}+5C_{\delta,k,\ell}\mathcal{E}(s;f)},$
    where $C_{\delta,k,\ell}$ is defined in Lemma 3.4;
@@
We conclude this section by commenting on the three terms appearing in the bounds in Theorem 3.5. The dependence on $4\sqrt{\ell s}\inf\limits_{p\in\mathbb{P}_{2r+1}}\|f(x)-p(x)\|_{L^{\infty}([\lambda_{\min},\lambda_{\max}])}$ captures the inherent cost of approximating a quadratic form with $f(\bm{A})$ using the Lanczos method, as in done in Line 3 of Algorithm 4. A similar term would also arise in an analysis of the Randomized SVD when matvecs are computed approximately using a Krylov subspace method. The $\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}$ term is due to the fact that our error can never be below the optimal rank $k$ approximation error. Finally, $\mathcal{E}(s;f)$ tells us that $\bm{Q}_{s}$ is a good orthonormal basis for low-rank approximation if there is a polynomial of degree at most s1s-1 that is large on the eigenvalues $\lambda_{1},\ldots,\lambda_{k}$ (which correspond to the top subspace of $f(\bm{A})$) and is small on the eigenvalues $\lambda_{k+1},\ldots,\lambda_{n}$.
As discussed in Section 3.2, one possible choice of polynomial would be an approximation to $f$, in which case $\mathcal{E}(s;f)$ would be close to $\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}$. In this case, ignoring the polynomial approximation term, we would recover a bound almost identical to the standard Randomized SVD error bound. In particular, when matvecs with $f(\bm{A})$ are implemented exactly, Randomized SVD can be shown to have expected error [rsvd]:
@@
When $f$ is a polynomial of degree at most $\max\{2r+1,s-1\}$, we can recover classical bounds for the randomized SVD. However, other possibly polynomials can also be chosen, so we might have $\mathcal{E}(s;f)\ll\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}$. In the next section, we give an example involving the matrix exponential to better illustrate why this is often the case.
@@ 3.5 Example bounds for the matrix exponential
In order to obtain explicit bounds for the approximation quality of Algorithm 4 one must obtain an explicit upper bound for $\mathcal{E}(s;f)$. This can be done, for example, by choosing a specific polynomial $p$ to obtain
$\mathcal{E}(s;f)\leq\|p(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}\max\limits_{i=1,\ldots,k}\left|\frac{f(\lambda_{i})}{p(\lambda_{i})}\right|.$
For monotone functions, a natural choice for $p$ is a scaled a shifted Chebyshev polynomial, as suggested in [MM15, tropp2023randomized]; this idea is also standard in the context of eigenvalue approximation [shmuel, paige1971computation, saad1980rates, underwood1975iterative].
One monotonic matrix function for which it is often desirable to obtain a low-rank approximation is the matrix exponential, $\exp(\bm{A})$. As highlighted in Remark Remark, this task arises in tasks ranging from network analysis [hpp], to quantum thermodynamics [chen_hallman_23, epperly2023xtrace], to solving PDEs [persson_kressner_23]. We emphasize that in many cases $\exp(\bm{A})$ admits an accurate low-rank approximation, even if $\bm{A}$ does not. Furthermore, at first glance one might think that one must require that the eigenvalues of $\bm{A}$ be negative for $\exp(\bm{A})$ to admit an accurate low-rank approximation. However, this is not the case. For any symmetric matrix $\bm{A}$ with eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}$, consider the shifted matrix $\bm{A}-\lambda_{1}\bm{I}$, whose eigenvalues are all non-positive. Since $\exp(\bm{A}-\lambda_{1}\bm{I})=\exp(-\lambda_{1})\exp(\bm{A})$, the matrix exponentials differ only by a scalar multiple. Consequently, any relative low-rank approximation error remains unchanged under shifts. Moreover, Algorithm 4 is unaffected by such a shift since $\mathcal{K}_{q}(\bm{A},\bm{\Omega})=\mathcal{K}_{q}(\bm{A}-\lambda_{1}\bm{I},\bm{\Omega})$.
In this section, we do a deeper dive into our main bounds for Krylov aware low-rank approximation when applied to the matrix exponential to better understand how they compare to bounds obtainable using the more naive approach of directly combining the randomized SVD with a black-box matrix-vector product approximation algorithm for $\exp(\bm{A})$. For conciseness of exposition, we focus on expectation bounds, and defer details of proofs to Appendix A.
We first demonstrates that, as discussed in the previous section, by choosing $p$ in $\mathcal{E}(s;\exp(x))$ to be a polynomial approximation to $\exp(x)$, we can recover the bounds of the randomized SVD [rsvd, Theorem 10.5], up to small factors accounting for the fact that $\exp(x)$ is not a polynomial.
@@
Note that as $\frac{\sqrt{\ell s}\gamma_{1,n}^{2r+2}}{2^{4r+1}(2r+2)!}\|\exp(\bm{A})\|_{2}\to 0$ and $\frac{\gamma_{1,n}^{s}}{s!}\to 0$, we recover the classical bound for the randomized SVD [rsvd, Theorem 10.5]. Furthermore, these two terms converge superexponentially to 0 as $s,r\to+\infty$. However, due to the $\frac{5k}{\ell-k-1}$, the above bound only implies a subexponential convergence to the error of the optimal rank $k$ approximation error, $\|\exp(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}$, in terms of the number of required matrix-vector products with $\bm{A}$, which scales as $\ell(r+s)$.
@@
Consider the setting of Theorem 3.5 with $f(x)=\exp(x)$. Define $\gamma_{i,j}=\lambda_{i}-\lambda_{j}$ to be the gap between the $i^{\text{th}}$ and $j^{\text{th}}$ largest eigenvalues. $\Gamma=\min\left\{1,2\frac{\gamma_{k,k+1}}{\gamma_{k+1,n}}\right\}$. Then, if $\ell-k\geq 2$ and $s\geq 2+\frac{\gamma_{1,k}}{\log\left(\frac{\gamma_{1,n}}{\gamma_{k,n}}\right)}$
@@
$\hskip 50.00008pt+\sqrt{1+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2^{-2(s-2)\sqrt{\Gamma}+3\gamma_{k,n}}}\frac{80k}{\ell-k-1}}\|\exp(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}.$
Consequently, if $s\geq 2+\frac{\gamma_{1,k}}{\log\left(\frac{\gamma_{1,n}}{\gamma_{k,n}}\right)}$ and
$\ell$ $=2k+1=O(k),$
$s$ $=O\left(\frac{\log(1/\varepsilon)+\gamma_{k,n}}{\sqrt{\Gamma}}\right),$
$r$ $=O\left(\gamma_{1,n}+\log(1/\tau)+\log\left(\frac{k(\log(1/\varepsilon)+\gamma_{k,n})}{\Gamma}\right)\right),$
for $\varepsilon,\tau\in(0,1)$, then we have
$\mathbb{E}\|\exp(\bm{A})-\textsf{{ALG}}_{k}(s,r;\exp(x))\|_{\textsf{{F}}}\leq\tau\|\exp(\bm{A})\|_{2}+(1+\varepsilon)\|\exp(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}.$
We believe that one can derive bounds that are independent of the eigenvalue gap $\Gamma$, reminiscent of the bounds in [MM15]. However, such bounds would likely require very different proof techniques.
4 Single-vector Krylov-aware low-rank approximation
In this section we consider a variant of Algorithm 4 that uses a single vector $\bm{\omega}\in\mathbb{R}^{n}$ as a starting block, rather than a block of vectors $\bm{\Omega}$. The resulting algorithm is given in Algorithm 5, and requires at most $k+s+r$ matvecs with $\bm{A}$ to compute an approximation to $f(\bm{A})$. This compares favourably to Algorithm 4, which requires $k(s+r)$ matvecs. The overall computational cost of Algorithm 5 is $k+s+r$ matvecs and $O(n(k+s+r)^{2})$ additional operations.
Algorithm 5 Single-vector Krylov-aware low-rank approximation
input: Symmetric $\bm{A}\in\mathbb{R}^{n\times n}$. Rank $k$. Matrix function $f:\mathbb{R}\to\mathbb{R}$. Degrees $s$ and $r$.
output: Low-rank approximation to $f(\bm{A})$: $\bm{Q}_{k+s}\bm{X}\bm{Q}_{k+s}^{\textsf{{T}}}$ or $\bm{Q}_{k+s}\llbracket\bm{X}\rrbracket_{k}\bm{Q}_{k+s}^{\textsf{{T}}}$
1:Sample a standard Gaussian random vector $\bm{\omega}$.
2:Run Algorithm 1 for $q=k+s+r$ iterations to obtain an orthonormal basis $\bm{Q}_{k+s}$ for $\mathcal{K}_{k+s}(\bm{A},\bm{\omega})$ and a tridiagonal matrix $\bm{T}_{q}$.
3:Compute $\bm{X}=f(\bm{T}_{q})_{1:d_{k+s,1},1:d_{k+s,1}}$ where $d_{k+s,1}=\dim(\mathcal{K}_{k+s}(\bm{A},\bm{\omega}))$. $\triangleright$ $\approx\bm{Q}_{k+s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{k+s}$
4:return $\widehat{\textsf{{ALG}}}(s,r;f)=\bm{Q}_{k+s}\bm{X}\bm{Q}_{k+s}^{\textsf{{T}}}$ or $\widehat{\textsf{{ALG}}}_{k}(s,r;f)=\bm{Q}_{k+s}\llbracket\bm{X}\rrbracket_{k}\bm{Q}_{k+s}^{\textsf{{T}}}$.
Many classical eigenvalue methods use Krylov subspace methods with small starting blocks, often as small as $\ell=1$ [shmuel, paige1971computation]. For instance, MATLAB’s eigs is based on a Krylov subspace initialized with a (random) starting block of size $\ell=1$, [^6] For reproducibility, MATLAB uses private random number stream. even when several eigenvalues are sought. In contrast, Krylov subspace methods for low-rank approximation have historically been analyzed for large block-sizes, i.e., $\ell\geq k$ [MM15, tropp2023randomized, drineasipsenkontopoulouismail, rsvd, gu_subspace]. The single-vector Lanczos method for low-rank approximation was recently analyzed in [MeyerMuscoMusco:2024]. Empirically, the single-vector Lanczos method often produces more accurate low-rank approximations using the same number of matvecs, though the block-Lanczos method comes with other advantages, such as being easier to parallelize. Recent work has also studied intermediate block-sizes $1\leq\ell\leq k$ in the context of low-rank approximation, which can balance accuracy and efficiency [moderateblocksize]. In what follows, we will focus on the case when the block-size is $\ell=1$, but we suspect that the analysis can be extended to any block-size.
A key observation made in [MeyerMuscoMusco:2024] is that the single-vector Lanczos method can be viewed as a block-Lanczos method with a highly structured starting block. Specifically, if $\bm{\omega}\in\mathbb{R}^{n}$ is a starting vector, then by Lemma 2.2 we have
$\mathcal{K}_{s+k}(\bm{A},\bm{\omega})=\mathcal{K}_{s}(\bm{A},\widehat{\bm{\Omega}})\quad\text{ where }\quad\widehat{\bm{\Omega}}:=\begin{bmatrix}\bm{\omega}&\bm{A}\bm{\omega}&\cdots&\bm{A}^{k-1}\bm{\omega}\end{bmatrix}$.
The benefit of this observation was that the existing analysis for the block-Lanczos algorithm for low-rank approximation could be used to analyze the single-vector Lanczos method for low-rank approximation, provided one could analyze the behavior of the structured starting block $\widehat{\bm{\Omega}}$. To see this, note that running Algorithm 5 is equivalent to running Algorithm 4 with $\widehat{\bm{\Omega}}$ instead of $\bm{\Omega}$. Hence, if we define
$\widehat{\bm{\Omega}}_{k}:=\bm{U}_{k}^{\textsf{{T}}}\widehat{\bm{\Omega}},\quad\widehat{\bm{\Omega}}_{n\setminus k}:=\bm{U}_{n\setminus k}^{\textsf{{T}}}\widehat{\bm{\Omega}}$,
then, assuming $\operatorname{rank}(\widehat{\bm{\Omega}}_{k})=k$ and applying Theorem 3.1, Lemma 3.3, [^7] Recall that Lemma 3.3 does not assume anything about the randomness of the starting block. and Lemma 2.3 we obtain the following structural bound
$\begin{split}\|f(\bm{A})-\widehat{\textsf{{ALG}}}_{k}(s,r;f)\|_{\textsf{{F}}}\leq&4\sqrt{k+s}\inf\limits_{p\in\mathbb{P}_{2r+1}}\|f(x)-p(x)\|_{L^{\infty}([\lambda_{\min},\lambda_{\max}])}+\\ &\sqrt{\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}+5\min\limits_{p\in\mathbb{P}_{s-1}}\left[\|p(\bm{\Lambda}_{n\setminus k})\widehat{\bm{\Omega}}_{n\setminus k}\widehat{\bm{\Omega}}_{k}^{-1}\|_{\textsf{{F}}}^{2}\max\limits_{i=1,\ldots,k}\left|\frac{f(\lambda_{i})}{p(\lambda_{i})}\right|^{2}\right]}.\end{split}$ (18)
Bounds for $\widehat{\textsf{{ALG}}}(s,r;f)$ can be obtained analogously. To obtain a probabilistic bound, it remains to control $\|p(\bm{\Lambda}_{n\setminus k})\widehat{\bm{\Omega}}_{n\setminus k}\widehat{\bm{\Omega}}_{k}^{-1}\|_{\textsf{{F}}}^{2}$ for a fixed polynomial $p$, which can be done using the results from [MeyerMuscoMusco:2024]. In particular, we have the following result, whose proof is deferred to Appendix B. [^8] Unfortunately, the proof technique cannot be used to prove expectation bounds, since the proof relies on bounding the inverse of a $\chi_{1}^{2}$ random variable, whose expectation does not exist.
Theorem 4.1.
Consider $\bm{A}\in\mathbb{R}^{n\times n}$ as defined in (1) with smallest and largest eigenvalues $\lambda_{\min}$ and $\lambda_{\max}$ respectively. Define $\gamma_{\min}=\min\limits_{\begin{subarray}{c}i,j=1,\ldots k\\ i\neq j\end{subarray}}\frac{|\lambda_{j}-\lambda_{i}|}{|\lambda_{\max}-\lambda_{\min}|}$. Then, with $\mathcal{E}(s;f)$ as defined in Eq. 16 the following inequality holds with probability at least $1-\delta$
$\|f(\bm{A})-\widehat{\textsf{{ALG}}_{k}}(s,r;f)\|_{\textsf{{F}}}\leq$ $4\sqrt{k+s}\inf\limits_{p\in\mathbb{P}_{2r+1}}\|f(x)-p(x)\|_{L^{\infty}([\lambda_{\min},\lambda_{\max}])}+$
@@ 4 Single-vector Krylov-aware low-rank approximation
$\sqrt{\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}+\widehat{C}_{\delta}\frac{k^{4}}{\gamma_{\min}^{2(k-1)}}\mathcal{E}(s;f)},$
where $\widehat{C}_{\delta}=\frac{2\pi(1+6\log(2/\delta))}{\delta^{2}}$.
We highlight the dependence on $\frac{1}{\gamma_{\min}^{2(k-1)}}$, which can be extremely large of $\bm{\Lambda}_{k}$ has tightly clustered eigenvalues. This term appears because of the dependence of the inverse of $\widehat{\bm{\Omega}}_{k}$ in (18), which becomes exponentially ill-conditioned in $k$ and depends severely on the separation of the eigenvalues in $\bm{\Lambda}_{k}$. In particular, if $\bm{A}$ has repeated eigenvalues, $\widehat{\bm{\Omega}}$ may become rank deficient. To illustrate, if $\bm{A}=\bm{I}$ then $\dim(\mathcal{K}_{s+k}(\bm{A},\bm{\omega}))=1$, independently of $s+k$. In particular, we have $\operatorname{rank}(\widehat{\bm{\Omega}})=1$ and $\gamma_{\min}=0$. Hence, the single-vector Lanczos method may fail to produce non-trivial low-rank approximations in degenerate cases when $\bm{A}$ has repeated eigenvalues.
However, as argued in [MeyerMuscoMusco:2024], provided $\gamma_{\min}\neq 0$, the dependence on $\gamma_{\min}$ is mild since $\mathcal{E}(s;f)$ usually decays rapidly in $s$. To illustrate, consider the case when $f(x)=\exp(x)$. In the proof of Corollary 3.7 we showed that
$\mathcal{E}(s;f)=O\left(2^{-2(s-2)\sqrt{\Gamma}+3\gamma_{k,n}}\|\exp(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}\right).$
Thus, to ensure $\widehat{C}_{\delta}\frac{k^{4}}{\gamma_{\min}^{2(k-1)}}\mathcal{E}(s;f)\leq\varepsilon\|\exp(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}$ it is sufficient set
$s=O\left(\frac{\log(1/\varepsilon)+\gamma_{k,n}+k\log(1/\gamma_{\min})+\log(\widehat{C}_{\delta})}{\sqrt{\Gamma}}\right).$
Hence, the dependence on $\frac{1}{\gamma_{\min}}$ only appears in a logarithmic factor. As we will see in Section 5, Algorithm 5 usually produce significantly more accurate low-rank approximations using a fixed computational cost.